OpenStudy (anonymous):

If b^2*c^2+b^2+c^2=243 a^2*c^2+a^2+c^2=1024 b^2*a^2+b^2+a^2=293 then what is (a+b+c)....?

OpenStudy (anonymous):

i want any idea to start....

OpenStudy (jack1):

substitution... lots of times... until you have = a ...?, b = ...? and c = ...?

OpenStudy (anonymous):

it kills lot of time...

OpenStudy (jack1):

b^2*c^2+b^2+c^2=243 therefore c^2 = 243 - b^2*c^2 - b^2 so c = sqrt (243 - b^2*c^2 - b^2) theres your start, sub into eqn 2 rearrange eqn 2 for a = ... something... sub THAT into eqn 3 rearrange so u end up with b = ...?

OpenStudy (anonymous):

hey i have a better option

OpenStudy (anonymous):

it is an interview question.... if we do like this... it takes 1 hour to calculate...

OpenStudy (anonymous):

ya tell me...@Aditya9826042721

OpenStudy (anonymous):

we do it step by step

OpenStudy (anonymous):

kk tell me your idea....

OpenStudy (anonymous):

first let a^2=x b^2 =y c^2=z

OpenStudy (anonymous):

sorry i have to go bcoz electricity gone i am on inverter

OpenStudy (anonymous):

if u there then rply to continue

OpenStudy (anonymous):

would u get the answer

OpenStudy (jack1):

im here ... shoot

OpenStudy (anonymous):

from first equation b^2c^2+b^2+c^2=243

OpenStudy (anonymous):

take b^2 out and add 1both the side

OpenStudy (anonymous):

as we get b^2(c^2+1)+c^2+1=243 +1

OpenStudy (anonymous):

now take out (b^2+1)(c^2+1)=244

OpenStudy (anonymous):

now similarly for all other equation (a^2+1)(c^2+1)=1025

OpenStudy (anonymous):

would u get it or not

OpenStudy (anonymous):

are any one there....? :p

OpenStudy (jack1):

yeah dude, i follow so far

OpenStudy (jack1):

youre breakin them down to quadratics, then factoring

OpenStudy (anonymous):

now we get three equation as about in multiplication form

OpenStudy (anonymous):

now multiply all three equation to get ((b^2+1)^2)((a^2+1)^2)((c^2+1)^2)=244*1025*294

OpenStudy (anonymous):

ya i am hear... Net problem sorry...

OpenStudy (anonymous):

now from either of equation like (b^2+1)(c^2+1)=244 put in last equation as ((244)^2)((a^2+1)^2)=244*1025*294 here u get equation of a solve and get a similarly for other

OpenStudy (anonymous):

hope u get it

OpenStudy (jack1):

nice

OpenStudy (anonymous):

thanks...........: )

OpenStudy (alekos):

been watching this. very clever!

OpenStudy (anonymous):

thanks a lot ..........